OpenStudy (anonymous):

The number of baby deliveries in a small hospital was sampled over 6 hour periods. Below is the sample: 11 12 7 16 4 8 9 7 3 4 9 4 6 12 9 8 4 6 12 From this data set, compute the sample mean, median, mode and midrange. Compute the 1st quartile and 3rd and interpret their meanings.

OpenStudy (anonymous):

do u know definition of mean median and mode?

OpenStudy (anonymous):

http://www.purplemath.com/modules/meanmode.htm

OpenStudy (anonymous):

mean is average so add em and divide by number of data entered

OpenStudy (anonymous):

median is the middle value between highest and lowest value

OpenStudy (anonymous):

mode is the most frequent number that occurs in the list so just look at it

OpenStudy (lalaly):

3 4 4 4 4 6 6 7 7 8 8 9 9 9 11 12 12 12 16 mean = 151/19 median = (19+1)/2 or tenth term mode is the mst frequent =4 midrange (16-3)/2 = 7.5 1st quartile = (19+1)/4 or the 5th term 3rd quartile = 9

OpenStudy (anonymous):

11+12 +7+16+4+8+9+7+3+4+9+4+6+12+9+8+4+6+12=151 total 19 values Mean = 151/19 =7.94736842 Arranging in ascending order 3,4,4,4,4,6,6,7,7,8,8,9,9,9,11,12,12,12,16. Total 19 values. Median is the 10th value = 8 Mode is the data that appears maximum number of times = 4 Median 8 divides the data into two halves. Median of 3,4,4,4,4,6,6,7,7,8 is the 5th value = 4 is the 1st quartile. Median of 8,9,9,9,11,12,12,12,16 is the 5th value = 11 is the third quartile

OpenStudy (anonymous):

thanks evetone!!! :)!

OpenStudy (lalaly):

:)

OpenStudy (lalaly):

do u want step by step?

OpenStudy (lalaly):

first Arrangethe numbers in ascending order 3 4 4 4 4 6 6 7 7 8 8 9 9 9 11 12 12 12 16

OpenStudy (anonymous):

yes step by step please!

OpenStudy (lalaly):

u hav 19 numbers now for the mean, u add up all the numbers and divide it by hw many numbers ther are so 11+12 +7+16+4+8+9+7+3+4+9+4+6+12+9+8+4+6+12=151 so the mean is 151/19 = 7.947

OpenStudy (lalaly):

the median is the middle value (19+1)/2 = 10th value, which is 8 Mode is the most frequent number which = 4 the mid range is the (largest value-lowest value)/2 = (16-3)/2=7.5 1st quartile = (19+1)/4 =the 5th term 3rd quartile = 3/4 ( 19+1) =15th term